HETAV PATEL

Grade 11,

Let a,b,c are in A.P. if p is the A.M of a and b and q is A.M of b and c.Then prove that b is the A.M between p and q

Let a,b,c are in A.P. if p is the A.M of a and b and q is A.M of b and c.Then prove that b is the A.M between p and q

Grade:11

1 Answers

Arun
6 years ago
a,b,c are in A.P.
Hence
2b = a+c
Similarily
2p = a+b...........(1)
&
2q = b+c.........(2)
Add. Equation 1 &2
2p +2q = a+2b +c
2p +2q = 4b
Hence
p +q = 2b
Hence b is the A.M of p & q
Hence proved